OpenStudy (anonymous):

What is the slope of the line that passes through (-2, 3) and (4, -5)? b. What is the equation of the line with slope 4 and y-intercept at (0, -3)?

hero (hero):

The equation of a line is y = mx + b, where m = slope. Insert the points into the formula to get: 3 = -2m + b -5 = 4m + b Isolate and solve for b in both equations to get: b = 2m - 3 b = -4m -5 Set b = b to get 2m - 3 = -4m - 5 Place like terms on the same side to get: 4m + 2m = -5 + 3 Combine like terms to get: 6m = -2 Divide both sides by 6 to get m = -2/6 Reduce the fraction to get m = -1/3 Therefore, the slope of the line is m = -1/3

OpenStudy (anonymous):

this was a good explanation but theres are my answer choices

OpenStudy (anonymous):